f(x)=9x^4-25^2 how to find the zero of the function algebraically
Answer by stanbon(75887)

f(x)=9x^4-25^2 how to find the zero of the function algebraically
-------
solve (3x^2-25)(3x^2+25) = 0
x^2 = 25/3 or x^2 = -25/3
-----
x = +/-5sqrt(3)/3 or x = +/-5sqrt(3)i/3
===============================================
Cheers,
Stan H.
====================